OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. x2 + 2x + y2 + 4y = 20

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

You would complete the squares (of x and y).

OpenStudy (anonymous):

Separating them into two equations?

OpenStudy (anteater):

Trying it again without the typos :) If you look at the x^2 + 2x and the y^2 + 4y, you can "complete the square". For example, x^2 + 2x are the first two terms of the perfect "square" x^2 + 2x + 1 or (x + 1)^2 and y^2 + 4y are the first two terms of the perfect "square" y^2 + 4y + 4 or (y + 2)^2

OpenStudy (mathmate):

\[x^2+2x+y^2+4y=20\] \[(x+1)^2+(y+2)^2=20+1+4\] I'll let you finish the problem.

OpenStudy (anteater):

Yes :)

OpenStudy (mathmate):

Yes, kind of, but they remain in the same equation (of the circle)

OpenStudy (anonymous):

Okay, let me finish the problem real quick! @anteater @mathmate

OpenStudy (anonymous):

Wait, but do I do anything to this? --> (x+1)2+(y+2)2 the other part equals to 25

OpenStudy (anonymous):

Do I distribute it, is what I mean.

OpenStudy (mathmate):

In a standard circle centred at h,k with radius r, the equation is \[(x-a)^2+(y-b)^2=r^2\] I'm sure you can figure out the centre and radius of the given circle.

OpenStudy (anonymous):

OH, now that actually helps. Thanks..I;m sure I can figure it out too.

OpenStudy (mathmate):

Sorry, the standard circle I gave had centre (a,b), and not (h,k) as I said. It was a typo. :(

OpenStudy (anonymous):

It's all good, :) It took me a good 2 minutes to figure that out haha :) Thanks! @mathmate and thanks @anteater

OpenStudy (anteater):

Welcome :) I figured I should just step aside since you guys were working away on it. :D

OpenStudy (anonymous):

Haha yeah thank you for that! I usually get confused when there's more than one person explaining at the same time..so sensible of you :))

OpenStudy (anteater):

Anteaters are known for sensibility. ;)

OpenStudy (anonymous):

But of course both of you helped in your own ways! (:

OpenStudy (anteater):

Sometimes.

OpenStudy (anonymous):

LOl haha ;))

OpenStudy (anteater):

Do you have any problems that involve ellipses or hyperbolas?

OpenStudy (mathmate):

It's been a pleasure. Catch up with you guys later!

OpenStudy (anteater):

Ok :)

OpenStudy (anonymous):

@anteater Nope! @mathmate ok :)

OpenStudy (anteater):

Ok :) I hope you have a great afternoon! :D

OpenStudy (anonymous):

Thanks, you too!